OpenStudy (anonymous):

the division property of equality could be used to solve which of the following equations? x^2+7x+6=0 10x=3 4-x=7 x/5=9

OpenStudy (anonymous):

10x = 3 i think you would divide both sides by 10 10x/10 = 3/10 x = 0.3

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no problem!